OpenStudy (anonymous):

PLEASE HELP!!! I have to rationalize the denominator of this expression, would square root 5 divided by square root 8x be square root 40x over 8x?

OpenStudy (anonymous):

Like would : \[\frac{ \sqrt{5}}{\sqrt{8x} }\] become: \[\frac{ \sqrt{40x} }{ 8x }\] ???

OpenStudy (ddcamp):

Yes, because \[\sqrt{5} * \sqrt{8x} = \sqrt{5*8x} = \sqrt{40x}\]

OpenStudy (anonymous):

okay, thank you so much @DDCamp!